# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 5.3. Duals of grid patches
For $x\in\mathsf{X}_{1}$, this follows from Lemma 5.5. For $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, we let $x_{k}=y({\boldsymbol{u}}_{k})^{*}\to x$, where $y({\boldsymbol{u}}_{k})\in Y_{1}$ (see Remark 4.7) can be chosen so that $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ and $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ for all $||{\boldsymbol{n}}||$ small. ∎
Definition 5.13.
For each Wang patch $w^{a}$, $a\in{\mathcal{A}}$ we define its symbol $\sigma(w^{j})=z_{0}\otimes z^{\prime}_{0}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single direction ${\boldsymbol{v}}_{2}^{\perp}$ pseudoline segment passes through $w^{j}$, we put $z^{\prime}_{0}=0$, but if two direction ${\boldsymbol{v}}_{2}^{\perp}$ pseudoline segments pass through $w^{j}$, we put $z^{\prime}_{0}=1$. Similarly, if a single direction ${\boldsymbol{v}}_{4}^{\perp}$ pseudoline segment passes through $w^{j}$, we put $z_{0}=0$, but if two direction ${\boldsymbol{v}}_{4}^{\perp}$ pseudoline segments pass through $w^{j}$, we put $z_{0}=1$.
Lemma 5.14.
Let $x\in\mathsf{X}$ be Penrose tiling covered by Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$. Then $z\otimes z^{\prime}=\{\sigma(w_{{\boldsymbol{n}}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence. Let $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$, and let ${\boldsymbol{u}}_{0}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$. If $x\in\mathsf{X}_{1}$ then $y({\boldsymbol{u}}_{0})\in Y_{1}$ and $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for some ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ let, $y({\boldsymbol{u}})=\varphi(x)$. Then $x$ is obtained from $T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}})^{*}$ for $||{\boldsymbol{t}}_{0}||<4/3$ by filling in a worm or cartwheel.
Proof.
First let $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$, so that $y({\boldsymbol{u}})\in Y_{1}$. By Definitions 5.3 and 5.13, $\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}})^{*})=\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}}))$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Thus $\{\sigma(w_{\boldsymbol{n}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence since $\{\sigma(g_{\boldsymbol{n}}({\boldsymbol{u}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is $\mathbb{Z}^{2}$ Sturmian by Lemma 5.4. For ${\boldsymbol{u}}_{0}$ as in the lemma, $y({\boldsymbol{u}}_{0})$ and $y({\boldsymbol{u}})$ are translates since they read the same Sturmian sequence. Thus there is a ${\boldsymbol{t}}_{0}$ so that $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$. For $y({\boldsymbol{u}})$ we have ${\boldsymbol{0}}\in R_{\boldsymbol{0}}$, whereas for $y({\boldsymbol{u}})^{*}$ the lower left vertex of $R_{\boldsymbol{0}}$ is ${\boldsymbol{0}}$. Thus $||{\boldsymbol{t}}_{0}||\leq||{\boldsymbol{f}}_{0}+{\boldsymbol{f}}_{1}||<4/3$, and (4.4) implies $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ and $y({\boldsymbol{u}})=\varphi(x)$, then as above, $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$. Fill the worm or cartwheel in $y({\boldsymbol{u}}_{0})^{*}$ the same way as $y({\boldsymbol{u}})^{*}$ is filled to get $x$. ∎
@@
The proofs of these Propositions are similar, but we separate them for conceptual clarity.
@@
For any $x\in\mathsf{X}$ and ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, there is $r^{\prime}_{1}>0$ with the following property: Let $w_{\boldsymbol{n}}(x)$ and $w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}(x)$ be adjacent Wang patches, with $\eta^{\prime}_{n_{0}}(x)\subseteq\eta_{n_{0}}$ the edge curve segment from ${\boldsymbol{d}}_{\boldsymbol{n}}$ to ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$. Then there exists ${\boldsymbol{s}}\in\eta^{\prime}_{n_{0}}(x)$ so that $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}})$. This follows from the FLC property for $\mathsf{X}$. Similarly, there is $r^{\prime\prime}_{1}>0$ so that $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}(x)\subseteq B_{r^{\prime\prime}_{1}}({\boldsymbol{s}})$ for some ${\boldsymbol{s}}\in\eta^{\prime}_{n_{1}}$. Let $r_{1}>\max\{r^{\prime}_{1},r^{\prime\prime}_{1}\}$.
Now fix $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$. Let $n_{0}\in\mathbb{Z}$ and ${\boldsymbol{s}}\in\eta_{n_{0}}(x)$. Then $B_{r_{1}}({\boldsymbol{s}})$ contains at least one Wang patch from the row $\{w_{(n_{0},n_{1})}(x)\}_{n_{1}\in\mathbb{Z}}$. In other words,
$\forall n_{0}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}\in\eta_{n_{0}}(x),\exists n_{1}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}).$ (5.3)
Similarly,
$\forall n_{1}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}\in\eta_{n_{1}}(x),\exists n_{0}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}).$ (5.4)
@@
Since ${\mathcal{V}}$ is not perpendicular to ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$, and since all the 00 and $1$-grid lines in $y({\boldsymbol{u}})$ are perpendicular to either ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$, ${\mathcal{V}}$ crosses every grid line $\ell_{1,n_{0}}(u_{1})$, $n_{0}\in\mathbb{Z}$, and every grid line $\ell_{0,n_{1}}(u_{0})$, $n_{1}\in\mathbb{Z}$. It follows from Lemma 5.12 that ${\mathcal{V}}$ crosses every edge curve $\eta_{n_{0}}(x)$, $n_{0}\in\mathbb{Z}$, and every edge curve $\eta_{n_{1}}(x)$, $n_{1}\in\mathbb{Z}$.
For a given $n_{0}\in\mathbb{Z}$, use (5.3) to find ${\boldsymbol{s}}\in{\mathcal{V}}\cap\eta_{n_{0}}(x)$ and $n_{1}\in\mathbb{Z}$ so that $w_{(n_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{1}}$. Similarly, use (5.4) to find ${\boldsymbol{s}}\in{\mathcal{V}}\cap\eta_{n_{1}}(x)$ and $n_{0}\in\mathbb{Z}$ so that $w_{(n_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{1}}$. Let =${\mathcal{L}}=\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:w_{{\boldsymbol{n}}}(x)\subseteq{\mathcal{V}}^{r_{1}}\}$. Since ${\mathcal{V}}^{r_{1}}=\bigcup_{{\boldsymbol{s}}\in{\mathcal{V}}}B_{r_{1}}({\boldsymbol{s}})$, it follows that ${\mathcal{L}}$ is thick in the sense of Definition 2.6. By Lemma 5.14, $\{\sigma(w_{\boldsymbol{n}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence, and by Lemma 2.7, it is determined by its values on the thick set ${\mathcal{L}}$. Using Lemma 5.14 again, $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for ${\boldsymbol{u}}_{0}=(0,0,u_{2},u_{3},u_{4})$, where $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$ and $u_{3}=\{-(u_{2}+u_{4})\}$. Since $x=y({\boldsymbol{u}})^{*}=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$, Lemma 4.5 implies ${\boldsymbol{u}}=K^{{\boldsymbol{t}}_{0}}{\boldsymbol{u}}_{0}$, so $x=y(K^{{\boldsymbol{t}}_{0}}{\boldsymbol{u}}_{0})^{*}$. ∎
@@
Case (b) is the case that $x$ has a cartwheel. There are two subcases: (i) $B_{r_{2}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{2}}$ contains an unfilled decagon in $y({\boldsymbol{u}})^{*}$. This decagon is filled in $x$, and that filling determines the direction of filling for all of the spokes. (ii) ${\mathcal{V}}$ crosses all five spokes, $l_{j}$, $j=0,1,\dots,4$. If ${\boldsymbol{s}}_{j}$ is the intersection of $l_{j}$ with ${\mathcal{V}}$, then there is a hexagon $H^{j}\subseteq B_{r_{2}}({\boldsymbol{s}}_{j})\subseteq{\mathcal{V}}^{r_{2}}$ in $y({\boldsymbol{u}})^{*}$. The filling of each $H^{j}$ determines the filling of the spoke it is in, which, in turn, determines the filling of the decagon. ∎

### v2

@@ 5.3. Duals of grid patches
First let $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$, so that $y({\boldsymbol{u}})\in Y_{1}$. By Definitions 5.3 and 5.13, $\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}})^{*})=\sigma(g_{{\boldsymbol{n}}}({\boldsymbol{u}}))$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Thus $\{\sigma(w_{\boldsymbol{n}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence since $\{\sigma(g_{\boldsymbol{n}}({\boldsymbol{u}}))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is $\mathbb{Z}^{2}$ Sturmian by Lemma 5.4. For ${\boldsymbol{u}}_{0}$ as described in the lemma, $y({\boldsymbol{u}}_{0})$ and $y({\boldsymbol{u}})$ are translates since they read the same Sturmian sequence. Thus there is a unique ${\boldsymbol{t}}_{0}$ so that $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$. For $y({\boldsymbol{u}})$ we have ${\boldsymbol{0}}\in R_{\boldsymbol{0}}$, whereas for $y({\boldsymbol{u}})^{*}$ the lower left vertex of $R_{\boldsymbol{0}}$ is ${\boldsymbol{0}}$. Thus $||{\boldsymbol{t}}_{0}||\leq||{\boldsymbol{f}}_{0}+{\boldsymbol{f}}_{1}||<4/3$, and (4.4) implies $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$.
If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ let $x_{n}\to x$ in $\mathsf{X}$. Let $(z\otimes z^{\prime})_{n}$ be the $\mathbb{Z}^{2}$ Sturmian sequence for $x_{n}$, and let ${\boldsymbol{u}}_{0,n}$ be the value of ${\boldsymbol{u}}_{0}$ (Lemma 5.4) for $(z\otimes z^{\prime})_{n}$. Then the definition of the tiling metric $\rho$ provides that $x_{n}\to x$ implies $(z\otimes z^{\prime})_{n}$ converges to a $\mathbb{Z}^{2}$ Sturmian sequence $z\otimes z^{\prime}$, which is the sequence read from $x$, (i.e., $z_{n_{0}}\otimes z^{\prime}_{n_{1}}=\sigma(w_{(n_{0},n_{1})}(x))$. If ${\boldsymbol{u}}=y^{-1}(\varphi(x))$ then since $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, $y({\boldsymbol{u}})^{*}$ has either an unfilled worm or an unfilled cartwheel. But also $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0,n}}y({\boldsymbol{u}}_{0,n})$ for some ${\boldsymbol{t}}_{0,n}$ with $||{\boldsymbol{t}}_{0,n}||<4/3$, and in fact ${\boldsymbol{t}}_{0,n}={\boldsymbol{t}}_{0}$ for all sufficently large $n$. So $y({\boldsymbol{u}})=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})$, but also $T^{{\boldsymbol{t}}_{n}}y({\boldsymbol{u}}_{0,n})^{*}=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0,n})^{*}\to x$. Thus $x$ is obtained from $y({\boldsymbol{u}})^{*}$ by filling in a worm or cartwheel. ∎
@@
The proofs of these Propositions are similar, but we separate them for conceptual clarity. We begin with the following two lemmas.
Lemma 5.17.
There exists an $r_{0}>0$ so that the following holds. Suppose $x\in\mathsf{X}$ with $x=T^{{\boldsymbol{t}}_{0}}x^{\prime}$ for $||{\boldsymbol{t}}_{0}||<4/3$. If $x[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$ for $||{\boldsymbol{t}}^{\prime}_{0}||<4/3$ then ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$.
This is essentially the locally free property for $\mathsf{X}$, but on a new scale.
Proof.
For the unmarked Penrose prototiles $p_{5}$ we compute $\Delta_{p_{5}}=2/5$, (see Definition 3.2). For any prototile set $p$, $\Delta_{\gamma p}=\gamma{\Delta_{p}}$, so $\Delta_{\gamma^{3}p_{5}}>4/3$. Consider the MLD topological conjugacy $(C^{-1})^{3}:\mathsf{X}\to\gamma^{3}\mathsf{X}$ where $\gamma^{3}\mathsf{X}\subseteq X_{\gamma^{3}p_{5}}$. Let $r_{0}>0$ be the radius for the local map $(C^{-1})^{3}$, (see Proposition 3.5). Now assume $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]=x[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$. Let $x_{3}:=(C^{-1})^{3}x\in\gamma^{3}\mathsf{X}$. By conjugacy, $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime}_{3}))[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime}_{3}))[B_{r_{0}}({\boldsymbol{0}})]$. Since $||{\boldsymbol{t}}_{0}||,||{\boldsymbol{t}}^{\prime}_{0}||<4/3<\Delta_{\gamma^{3}p}$, Lemma 3.3(1) shows ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$. ∎
Lemma 5.18.
There is an $r^{\prime}_{1}>0$ so that for all $x\in\mathsf{X}_{1}$:
$\forall n_{0}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}_{0}\in\lambda_{1,n_{0}},\exists n^{\prime}_{1}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{0}),$ (5.5)
and
$\forall n_{1}\in\mathbb{Z}\text{\rm\ and }{\boldsymbol{s}}\in\lambda_{0,n_{1}},\exists n^{\prime}_{0}\in\mathbb{Z}\text{\rm\ so\ that }w_{(n_{0},n_{1})}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{)}.$ (5.6)
Proof.
There is an $r^{\prime}_{1}>0$ so that for any ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$ the following holds: If ${\boldsymbol{s}}_{0}\in\lambda^{\prime}_{0,n_{1}}=\overline{{\boldsymbol{d}}_{\boldsymbol{n}}\,{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}}$ and ${\boldsymbol{s}}_{1}\in\lambda^{\prime}_{1,n_{0}}=\overline{{\boldsymbol{d}}_{\boldsymbol{n}}\,{\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}}$, then $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{0})$ and $w_{\boldsymbol{n}}(x)\cup w_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}(x)\subseteq B_{r^{\prime}_{1}}({\boldsymbol{s}}_{1})$. Thus for any $n_{0}\in\mathbb{Z}$ and ${\boldsymbol{s}}\in\lambda_{1,n_{0}}(x)$, $B_{r^{\prime}_{1}}({\boldsymbol{s}})$ contains at least one Wang patch from the row $\{w_{(n_{0},n_{1})}(x)\}_{n_{1}\in\mathbb{Z}}$. This proves (5.5), and the proof of (5.6) is the same. ∎
@@
For each ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, let $z_{n_{0}}\otimes z^{\prime}_{n_{1}}:=\sigma(w_{\boldsymbol{n}}(x))$. By the assumption that $x\in\mathsf{X}_{1}$, Lemma 5.14 implies that $z\otimes z^{\prime}\in Z\otimes Z$ is a $\mathbb{Z}^{2}$ Sturmian sequence, Moreover, Lemma 5.14 implies that if $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$ and $u_{3}=\{-(u_{2}+u_{4})\}$, then for ${\boldsymbol{u}}_{0}=(0,0,u_{2},u_{3},u_{4})$ we have $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for $y({\boldsymbol{u}}_{0})\in Y_{1}$ and (unique) ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$.
Let $r_{1}\geq\max\{r_{0},r^{\prime}_{1}\}$, (Lemmas 5.17 and 5.18). The proof now proceeds by showing that we can find $z\otimes z^{\prime}$ and ${\boldsymbol{t}}_{0}$ from $x[{\mathcal{V}}^{r_{1}}]$.
@@
Since ${\mathcal{V}}$ is not perpendicular to ${\boldsymbol{v}}_{0}$ or ${\boldsymbol{v}}_{1}$, it crosses every 00-grid line $\ell_{0,n_{1}}$ and every $1$-grid line $\ell_{1,n_{0}}$ in $y({\boldsymbol{u}})$. Thus by Lemma 5.11, ${\mathcal{V}}$ crosses every bent 00-grid line $\lambda_{0,n_{1}}$ and every bent $1$-grid line $\lambda_{1,n_{0}}$ in $x$.
For each $n_{0}\in\mathbb{Z}$, we use (5.5) to find ${\boldsymbol{s}}_{0}\in{\mathcal{V}}\cap\lambda_{1,n_{0}}$ and $n^{\prime}_{1}\in\mathbb{Z}$ so that $w_{(n_{0},n^{\prime}_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}_{0})\subseteq{\mathcal{V}}^{r_{1}}$. Similarly, we use (5.6) to find ${\boldsymbol{s}}_{1}\in{\mathcal{V}}\cap\lambda_{0,n_{1}}(x)$ and $n^{\prime}_{0}\in\mathbb{Z}$ so that $w_{(n^{\prime}_{0},n_{1})}(x)\subseteq B_{r_{1}}({\boldsymbol{s}}_{1})\subseteq{\mathcal{V}}^{r_{1}}$. Let $J:=\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:w_{{\boldsymbol{n}}}(x)\subseteq{\mathcal{V}}^{r_{1}}\}$. Since ${\mathcal{V}}^{r_{1}}=\bigcup_{{\boldsymbol{s}}\in{\mathcal{V}}}B_{r_{1}}({\boldsymbol{s}})$, we see that $\pi_{0}(J)=\pi_{1}(J)=\mathbb{Z}$. Thus $x[{\mathcal{V}}^{r_{1}}]$ determines $(z\otimes z^{\prime})[J]$, which by Lemma 2.6 determines $z\otimes z^{\prime}$.
Now suppose $x^{\prime}=y({\boldsymbol{u}}_{0})^{*}$ from above, so that $x=T^{{\boldsymbol{t}}_{0}}x^{\prime}$ with $||{\boldsymbol{t}}_{0}||<4/3$. Let ${\boldsymbol{t}}^{\prime}_{0}\in\mathbb{R}^{2}$ satisfy $||{\boldsymbol{t}}^{\prime}_{0}||<4/3$ and suppose $x[{\mathcal{V}}^{r_{1}}]=(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[{\mathcal{V}}^{r_{1}}]$, (noting this holds for ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$). Then since $r_{1}\geq r_{0}$, we have $(T^{{\boldsymbol{t}}^{\prime}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]=(T^{{\boldsymbol{t}}_{0}}x^{\prime})[B_{r_{0}}({\boldsymbol{0}})]$ and Lemma 5.17 implies ${\boldsymbol{t}}^{\prime}_{0}={\boldsymbol{t}}_{0}$. ∎
@@
Case (b) is the case that $x$ has a cartwheel. There are two sub-cases: (i) $B_{r_{2}}({\boldsymbol{s}})\subseteq{\mathcal{V}}^{r_{2}}$ contains an unfilled decagon in $y({\boldsymbol{u}})^{*}$. This decagon is filled in $x$, and that filling determines the direction of filling for all of the spokes. (ii) ${\mathcal{V}}$ crosses all five spokes, $l_{j}$, $j=0,1,\dots,4$. If ${\boldsymbol{s}}_{j}$ is the intersection of $l_{j}$ with ${\mathcal{V}}$, then there is a hexagon $H^{j}\subseteq B_{r_{2}}({\boldsymbol{s}}_{j})\subseteq{\mathcal{V}}^{r_{2}}$ in $y({\boldsymbol{u}})^{*}$. The filling of each $H^{j}$ determines the filling of the spoke it is in, which, in turn, determines the filling of the decagon. ∎